SOHO DAILY REPORT Operations Day - 017/2001 (17 January 01) Mission Day 1874 SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC Ref. #: SH-FOT-DR-010117 ===================================================================== PASS #1 Madrid DSS66 0701-1432 PASS #2 Goldstone DSS16 1424-2250 ********************************************* PASS #1 Madrid DSS66 0701-1432 Anomalies: G01-0013 0734 Power Hit at IMOC caused loss of string 4 and IRTS 3, as well as ECS. NRT: 1254 - 1356 Activities Summary: - SSR Dump - Collect MDI-M (0840 - 0856, 1322 - 1356) - Collect MDI-H (1321 - 1400, 1406 - 1432) - Min/Max data collection - OBT Distribution to Swan/Celias - SSR to Record ------------------ PASS #2 Goldstone DSS16 1424-2250 Anomalies: S01-0007 1627 UVCS parameter UTTHDIT tripped yellow low limit S01-0008 2045 CDS not acknowledging OBDH blocks 2 S/W anomalies received. NRT: 1446 - 1528 1541 - 1618 1639 - 1742 1850 - 2045 2058 - 2229 Activities Summary: - Execute OCD #1450, UVCS and CDS substitution heater turn OFF - Uplink Delayed Loads (VIRA10118, NONCOHO_01017) - Add UVCS and LASCO to the experiment off-pointing flag - SSR To Record ------------------ -- End of Report --